Converting Scores 7. A T-score of 60 would equal a z-score of 8. An Intelligence score of 130 would equal a z-score of 9. A T-score of 30 would equal an Intelligence score of 10. A z-score of -3 would equal a T-score of 11. A person who scores in the sth stanine would fall in quartile 12. A score at the 50th percentile would fall into stanine 1. A score at the 20% percemtile would all into quartile 14. A score at the 5oth percentile would equal an Intelligence score of 15. A score at the 50th percentile would equal a t-score of
Answer #1

7)

T = z*10 + 50

So, z = (t-50)/10 = (60-50)/10 = 1

8)

IQ Score = 100 + 15*Z

So, Z = (IQ – 100)/15 = (130-100)/15 = 2

9)

When T = 30,

Z = (30-50)/10 = -2

So,

IQ Score = 100 – 2 *15 = 70

10)

When z = -3,

T = z*10 + 50 = 50 – 3*10 = 20

PS: We are only allowed to answer 4 parts per question